Circular No. 7558

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

SUPERNOVA 2001F IN IC 867 M. Modjaz and W. D. Li, University of California at Berkeley, on

behalf of LOTOSS (cf. IAUC 7514), report the discovery of an apparent supernova (mag ∼ 16.5) on an unfiltered image taken with the 0.8-m Katzman Automatic Imaging Telescope (KAIT) on Jan. 6.5 UT (confirmed on several images taken on Jan. 9.5 under poor weather conditions). The new object is located at α = 13^{h}17^{m}20^{s}.05, δ = +20^{o}38′09′′.8 (equinox 2000.0), which is 4′′.3 east and 7′′.7 south of the nucleus of IC 867. A KAIT image of the same field taken on 2000 June 20.2 showed nothing at this position (limiting mag ∼ 19.0).

COMET C/1999 T1 (McNAUGHT-HARTLEY) D. Schleicher, Lowell Observatory, reports that he obtained eight sets

of narrowband photometry of comet C/1999 T1 on 2000 Dec. 28 and 2001 Jan. 2 with the Hall 1.1-m telescope at Lowell Observatory, yielding the following averaged results: log Q(OH) = 28.67; log Q(CN) = 26.10; log Afρ = 3.06 (cf. IAUC 7342). The equivalent log Q(water; vectorial) is 28.76. No significant temporal or aperture variations were observed.

SUNGRAZING COMETS Further to IAUC 7548, D. Hammer reports his measurements for ad-

ditional Kreutz sungrazing comets found by J. Danaher (C/1998 B2), X. Leprette (Nov. comets), and M. Oates (Dec. comets) on SOHO website images taken with the C2 coronograph; C/1998 B2 was also visible in C3 data. The reduced observations and orbital elements by B. G. Marsden are given on the MPECs cited below.

Comet  1998 UT  α_{2000}  δ_{2000}  MPEC
C/1998 B2 Jan. 25.588  20^{h}48.^{m}0  −20^{o}20′  2001-A32
C/1998 V2 Nov. 7.309  14 44.0  −17 35  2001-A31
C/1998 V3  7.309  14 44.7  −17 37  2001-A31
C/1998 V4  9.121  14 50.8  −18 07  2001-A35
C/1998 W4  24.188  15 58.2  −22 11  2001-A31
C/1998 W5  24.692  16 01.4  −22 22  2001-A31
C/1998 W6  25.564  16 04.4  −22 30  2001-A35
C/1998 X9 Dec. 14.246  17 30.8  −25 03  2001-A32
C/1998 X10  14.271  17 30.3  −25 04  2001-A32
2001 January 9  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2001 CBAT  Daniel W. E. Green